can you please help me solve this:Ax+By=-28 contains the points (-8,4)(-11,9)
Answer by tutorcecilia(2152)

Ax+By=-28 contains the points (-8,4)(-11,9)
Find the slope of the line:
m=(y-y2)/(x-x2) [Use the formula for a slope]
m=(9-4)/(-11-(-8)=5/-3 [Plug-in the values]
.
(y-y2)=m(x-x2) [Use the point-slope equation of a line because you have a point and now you have the slope]
.
(y-9)=-5/3(x-(-11) [Plug-in the values]
y-9=-5/3x-55/3 [Simplify]
5/3x+y=-55/3+9 [Isolate the variables and the numbers to either side]
5/3x+y=-55/3+9 [Simplify]
5/3x+y=-28/3